Accelerator Update for April 21, 2008 - April 23, 2008
Monday April 21
The day shift began with Operations monitoring the Tevatron (TeV) store 6068 (75.04E30) and the Recycler stash (241.9E10), with the Antiproton Source (Pbar) stacking (30.9E10), with the Recycler, the Main Injector, the Booster and Linac operating normally, with the Collider Detector at Fermilab (CDF) and DZero (D0) using the colliding beams, and with the NuMI and SciBooNE / MiniBooNE (BNB) experiments taking beam.
At 8:45 AM, a Linac Klystron RF station (KRF3) tripped off. It reset.
A RF expert began repairing MIRF16 at 9 AM. He had completed his repairs within 30 minutes.
A Pbar expert reported at 9:20 AM that they were officially switching to the new rapid transfer aggregate, “Rec XFRs – One Shots,” for all future transfers of antiprotons to the Recycler.
Operations rebooted the Pbar font end (FE) at 9:26 AM per an expert’s request.
At 11:06 AM, Operations reported a problem with a Booster extraction kicker (B:MKS06). Experts investigated and replaced a charging box, but after the repairs it was still missing some pulses. Experts found and fixed some bad connectors.
At 11:15 AM, Operations reported a problem with a Booster power supply (B:NOKR). Experts investigated. They said it looked like someone had accidentally bumped into it and turned it off. The experts had it timed in and back on within 17 minutes.
At 11:48 AM, KRF6 tripped off. It reset.
At 12:25 PM, a Linac RF station (LRF1) tripped off. It reset.
Operations terminated store 6068 at 1:02 PM.
Water system experts made up water for the TeV LCW system at 1:17 PM.
Operations put the TeV into shot setup at 1:20 PM.
At 2:09 PM, KRF7 tripped off. It reset.
Water system experts made up water for the MI RF and Magnet LCW systems at 2:34 PM.
At 2:46 PM, a Recycler expert took out the new RF leveling board and replaced it with the old board. The new board wasn’t operating correctly.
A water system expert made up water for the MI-60 LCW system at 2:47 PM.
Operations established store 6069 at 2:55 PM with an initial luminosity of 250.04E30.
At 2:57 PM, Operations reported that the MIRF system was tripping on “APG Watchdog” trips. Operations contacted an expert who reported there was a spike on MIRF1. He reset everything.
At 3:38 PM, KRF2 tripped off on a LCW flow indication. It reset.
At 3:45 PM, KRF4 tripped off. It reset.
Operations reported at 3:56 PM that the Recycler had stashed antiprotons two times during the day shift.
At 4:01 PM, the P1 transfer line beam permit tripped off. It reset.
At 4:05 PM, the MIRF system tripped off. It reset.
At 5:26 PM, KRF6 tripped off. It reset.
At 8:07 PM, KRF3 tripped off. It reset.
At 8:08 PM, KRF1 tripped off. It reset.
At 9:15 PM, Operations reported MI LCW resistivity alarms. Operators investigated, but could not find any problems. Operations informed an expert. The resistivity was getting better.
Operations reported at midnight that the Recycler had stashed five times during the evening shift.
Tuesday April 22
The midnight shift began with Operations monitoring store 6069 (79.125E30) and the stash (211.7E10), with Pbar stacking (32.3E10), and with NuMI and BNB taking beam.
At 4:25 AM, KRF7 tripped off. It reset.
At 4:34 AM, the MIRF Watchdog tripped on a Cavity LCW Temperature indication. It reset.
Operations terminated store 6069 at 5:04 AM.
Operations put the TeV into shot setup at 5:15 AM.
Operations established store 6070 at 6:55 AM with an initial luminosity of 241.0E30.
Operations reported at 7:56 AM that the Recycler had stashed four times during the midnight shift.
A mechanical technician began repairs on the C4 air compressor at 8:36 AM. He completed the repair about an hour later.
At 9:37 AM, KRF7 tripped off. It reset.
A MI expert began a multibatch slipstacking study at 9:58 AM.
At 9:58 AM, a Booster vacuum pump (B:IPRF23) tripped off and caused BRF13 and 14 to trip off. Everything reset.
A Pbar expert reported at 11:31 AM that he was increasing the gradient for the lithium lens.
At 12:51 PM, KRF3 tripped off. It reset.
At 2:53 PM, KRF7 tripped off. It reset.
At 3:58 PM, BRF10 tripped off. It reset.
Operations reported at 4:02 PM that the Recycler had stashed five times during the day shift.
At 5:02 PM, KRF2 tripped off. It reset.
At 6:13 PM, KRF5 tripped off. It reset.
Operations terminated store 6070 at 10:06 PM.
Operations put the TeV into shot setup at 10:17 PM.
At 10:35 PM, Operations reported that a problem with the decelerate aggregate was delaying shot setup. Operations contacted a TeV expert who felt this was a controls problem.
A controls expert reported at 11:36 PM that he was reverting to a previous version of “CLIB;” it was conflicting with one of the TeV ACL scripts. Operations continued with shot setup.Operations reported at 11:57 PM that the Recycler had stashed four times during the evening shift.
Wednesday April 23
The midnight shift began with the TeV in shot setup, with Operations monitoring the stash (323.0E10), with Pbar stacking (86.0E10), and with NuMI and BNB taking beam.
At 12:44 AM, LRF2 tripped off. It reset.
Operations established store 6071 at 1:14 AM, but unfortunately the TeV suffered a sector F1 quench just after the store was established. Operations informed experts.
A TeV expert reported at 6:20 AM that the D17 separator was the possible cause for the F1 quench.Operations reported at 7:55 AM that the Recycler had stashed five times during the midnight shift.
The Plans for Wednesday
During the 9 o’clock meeting, two Recycler beam valves closed – the stash was lost. The Run Coordinator wants the machines to get back to stacking, stashing, and storing as soon as possible. He will allow Switchyard experts to conduct beam studies and after they have finished he may allow a MI heat exchanger flushes.
Comments and Suggestions
|last modified 04/23/2008 email Fermilab|
|Security, Privacy, Legal|